1) Nuthatch Corporation began its operations on September 1 of the current year. Budgeted sales for the first three months of business are $260,000, $350,000, and $400,000, respectively, for September, October, and November. The company expects to sell 30% of its merchandise for cash. Of sales on account, 80% are expected to be collected in the month of the sale and 20% in the month following the sale.

The cash collections in November from accounts receivable are:

2)Division A of Mocha Company has sales of $155,000, cost of goods sold of $83,000, operating expenses of $43,000, and invested assets of $150,000.

What is the rate of return on investment for Division A?
Answer #1

1. Calculation of cash collections in November from accounts receivable

given 30% are cash sales in any month balance 70% is credit sales

credit sales in November = 400000*70% = 280000

credit sales in December = 350000*70% = 245000

Particulars Calculation Amount in $
cash collections in the month of sale ( nov Credit sale * 80%)   280000*80% 224000
cash collections from Oct month (oct credit sale * 20%) 245000*20% 49000
Total cash collections from account receivables   273000

2. Calculation of rate of return for division A

Net profit = sales - costs of goods sold - operating expenses

= 155000-83000-43000

=29000

Return on investment = Net profit / investment *100

= 29000 /150000 *100

= 19.33%

Thank You!